I am having problems with solving these two
Solve: X^2=12
and solve: X^2-80=0
Thanks a million

Answer by checkley75(3666)

IS THIS FORM BETTER?????
x^2=12
x=sqrt12
x=2sqrt3
x=2*1.732
x=3.464 answer.
x^2=80
x=sqrt80
x=4sqrt5
x=4*2.236
x=8.944 answer.